OpenStudy (anonymous):

Help! Make a table and graph f(x)=1/2x^2

OpenStudy (anonymous):

what are you working on???

OpenStudy (anonymous):

Functions,

OpenStudy (anonymous):

oh okay you working on fuction tables okay working it out..

OpenStudy (anonymous):

I'm just not sure how to start it.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

?

OpenStudy (anonymous):

idk don't know it

OpenStudy (anonymous):

The table should have two columns, for x and y. Here we'll assume that y =f(x); 'y is a function of x'. So pick any random values of the variable x, plug them in the given function and find the value, it will give one point (x,y). With enough points you can simply mark them on the Cartesian plane and join for the graph of the function. \[y = \frac{1}{2x^2}\]For x = 1, \[ y = \frac{1}{2 \times 1^2} = \frac{1}{2}\]So the point is (1, 1/2) Similarly for x = 2, y = 1/8, thus the point is (2, 1/8) and so on.

OpenStudy (anonymous):

yea..

OpenStudy (anonymous):

Thats very confusing, so the outputs are going to be in fractions?

OpenStudy (anonymous):

yea..